In baseball, a player's batting average is the total number of hits divided by the total number of times at bat. A baseball player has 13 hits after 41 times at bat. How many additional consecutive times must the batter hit the ball to achieve a batting average of .440?


Answer by richwmiller(17219)

(13+x)/(41+x)=.44
x=9
check
(13+9)/(41+9)=.44
22/50=.44
ok